OpenStudy (anonymous):

When a polynomial f(x) is divided by x+1, the reminder is a. If f(2x+1) is divided by x+1, the reminder is...? (in terms of a) @Callisto again haha sorry for that

OpenStudy (anonymous):

is it a? f[2(-1)+1]=f(-1) same as f(x) ??

OpenStudy (callisto):

Hmmm... Not really.

OpenStudy (anonymous):

oh why?

OpenStudy (callisto):

Let me check something first...

OpenStudy (anonymous):

take you time xd

OpenStudy (callisto):

Haha! f(x) is divided by x+1, the reminder is a. => f(x) = (x+1) Q(x) + a Check: f(-1) = a f(2x+1) = [(2x+1) + 1] Q(2x+1) +a = (2x+2) Q(2x+1) + a = 2(x+1) Q(2x+1) + a So, f(-1) = ... XD Haha, you're right XD

OpenStudy (anonymous):

haha i know more about polynomials now xD thank you!!!!

OpenStudy (callisto):

You're welcome!!!